if a teller has 54 $5.00 and $20.00 bills in her drawer that total $780.00. How many of the bills are $5.00?
Answer by rapaljer(4671)

Let x = number of $5 bills
54-x = number of $20 bills

The equation is based upon the values of the bills times the number of bills:
5(x) + 20(54-x) = 780
5x + 1080 - 20x = 780
-15x + 1080= 780
-15x = 780-1080
-15x=-300
x=-300/-15= 20 $5 bills
54-x = 54-20 = 34 $20 bills

Check:
20*$5 + 34*$20
$100 + $680
$780

It checks!

For additonal explanation, examples, and exercises like these Word Problems, please see my own website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Look for "Basic, Intermediate, and College Algebra: One Step at a Time". Choose either "Basic Algebra" or "Intermediate Algebra" and look in Chapter 1 in both of these topics for "Word Problems". Many of these exercises are then solved on the "MATH IN LIVING COLOR" pages that are associated with these sections.

R^2

Dr. Robert J. Rapalje, Retired
Seminole Community College
Altamonte Springs, FL 32714